• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Fansler

    Log entry time 07:26:34 on June 18, 2010

    Entry number 327230

    keyword=Left Diple NMR

    The left Dipole NMR was locked up. I power cycled the APC switch and the NMR appears to be functioning normally. hareboot17 outlet 8.